# and quantum information science

### v1

@@ 3 Dialogue
The dialogue below is inspired by Galilei’s Dialogue Concerning the Two Chief World Systems and is similar to discussions one hears during conferences. We imagine a conversation among three participants: Sofia represents our view of complexity. Complexio acts as an interested audience member aiming to clarify the literature, albeit sometimes with technical questions. Sagredo mediates between the others.
Sagredo: Finally, we resolved to meet today to clarify the past decade’s central developments in complexity.
@@
Sofia: Complexity has diverse applications. One can use it to characterize quantum systems’ dynamics and thermalization, solve optimization problems, capture chaotic properties, describe quantum information scrambling…Shall I continue?
Sagredo: These applications sound interesting, but complexity seems intangible. In most cases, it is difficult or impossible to compute.
Sofia: Fair objection. Nevertheless, heuristic arguments reveal how a state’s quantum complexity changes under generic state evolution. Also, bounding complexity is easier than calculating it, and bounds suffice for many purposes.
Sagredo: Sofia, even finding bounds—other than the simple bounds known already—seems nontrivial. How can we make progress?
Sofia: At least two approaches can improve current bounds on complexity. One approach centers on random unitary circuits, which form a toy model for chaotic dynamics. Invoking random unitary circuits, we can apply differential topology and algebraic geometry to manifolds formed by unitaries. Also, when considering random unitary circuits, we can apply unitary designs—probability distributions that cannot be distinguished from the Haar measure in polynomial time.
The second approach centers on Nielsen’s notion of complexity. Suppose we want to perform some unitary on $n$ qubits. Each of many circuits implements that unitary. The unitary group $\mathrm{SU}(2^{n})$, equipped with a continuous norm, forms a differential manifold. Nielsen casts the optimal circuit as a path on the manifold. The geodesic’s length bounds the unitary’s quantum complexity [10, 14]. This bound is often imprecise, but it may be useful.
@@
Sofia: Certainly! Kolmogorov complexity is the length of the shortest computer program that produces a target output [15]. Another notion, Krylov complexity, quantifies operator growth—an operator’s spreading, under evolution in the Heisenberg picture, across the space of operators [16]. Similarly, spread complexity quantifies a state’s expansion across the state space under evolution [17].
@@
Sofia: Both have interesting physical applications and are relatively simple to calculate. Reference [16] introduced Krylov complexity to distinguish chaotic from integrable dynamics of finite-size quantum many-body systems that lack semiclassical limits. Spread complexity relates to spectral form factors, which characterize many-body quantum-systems’ energy spectra [17]. Moreover, to calculate the Krylov or spread complexity, one must diagonalize (or tridiagonalize) only the Hamiltonian’s projection onto a small subspace. Therefore, one can calculate these complexities relatively easily.
@@
Sagredo: I have heard that the holography duality is an equivalence between gravitational physics in a bulk geometry and a quantum theory defined on its boundary. Holographic duals usually simplify difficult quantum problems. Does quantum complexity have a holographic interpretation?
Sofia: We hope so! Quantum complexity has been proposed to have an Einstein-Rosen bridge’s (ERB’s) volume as a holographic dual. An ERB is a spacetime region that connects a maximally extended Schwarzschild black hole’s two boundaries. In the holographic context, the black hole lives in AdS spacetime and is dual to the thermofield-double state of a CFT on the spacetime’s boundary. The ERB’s volume was proposed to be a holographic dual to the boundary state’s quantum complexity [18, 19]. More-complicated geometric quantities were conjectured to be dual to quantum complexity, too [20, 21, 22, 23, 24].
@@
Sofia: Yes. The geometric quantities behave similarly to the complexity of a quantum state evolving under random dynamics. One important behavior is linear growth until late times, achievable by geometric quantities in a semiclassical gravitational setting. Another behavior is the switchback effect, in which a perturbation delays growth.
More evidence emerges from a class of two-dimensional bulk models. There, a notion of wormhole length was formulated nonperturbatively. This length saturates at times exponentially long in the black hole’s entropy [25]. This behavior mirrors the saturation of complexity, under random dynamics, at times exponentially long in the system size. Moreover, in [26, 27], the authors show that a state’s spread complexity parallels a wormhole’s volume, at least in a two-dimensional gravitational setting.
@@
Sofia: The map I just mentioned—between a wormhole’s volume and spread complexity—hints that spread complexity is appropriate. Yet Refs. [28, 29, 30, 31, 32, 33, 34] point to other notions of complexity. We are not yet certain.
@@
Complexio: The community sounds awfully uncertain about complexity in holographic contexts. Why bother working on the topic?
@@
Sofia: First, the gravitational and quantum settings would merit analysis even in the absence of any duality. We want to understand black holes’ interiors and to optimize quantum circuits. Complexity has facilitated these goals. Second, we do hope to replace the many conjectures with a smaller number of theorems.
@@
Complexio: The community seems to me to be far from theorems. To support a bulk-boundary map for Nielsen’s complexity, people constructed toy examples involving free QFTs. Yet free QFTs are not known to be dual to any gravitational systems. Rather, certain strongly coupled QFTs are. So how can the toy examples support the map?
@@
Sofia: We expect some qualitative properties of QFTs not to depend on the QFT’s coupling strength. An example is the structure of the UV divergence of a QFT state’s complexity.222 The complexity contains multiple terms, which scale differently with the UV cutoff. Therefore, some people call “the structure of the UV divergence” “the UV divergences.” The UV divergence of the entanglement entropy offers another example.
@@
Sagredo: Indeed, entanglement entropy provides a striking example of a holographic mapping that passed non-trivial consistency checks. The dual to a boundary QFT state’s entanglement entropy is the area of a minimal surface in the geometry dual to that state. The dual objects have been identified clearly in this example for various states.
@@
Complexio: In contrast, according to the “complexity=anything” prescription, complexity parallels each of infinitely many gravitational quantities [23, 24, 35]. Every such quantity exhibits linear growth and the switchback effect. I can pick any conjecture I like, and all of them define a geometrical dual to quantum complexity. Does this multiplicity make sense? Is the cornucopia of holographic conjectures useful?
@@
Sofia: The multiplicity does offer an upshot. As we discussed, quantum state complexity depends on a gate set and a reference state. We are free to choose each. This freedom mirrors the ambiguity in the complexity’s gravitational dual.
Still, I understand Complexio’s concern: some symmetries or other properties could restrict the gravitational dual to one geometric object, and we could be blind to those properties. One might overcome such blindness, or confirm a lack thereof, as follows. First, show that the complexity has exactly as many degrees of freedom as the gravitational analogue. Building on this result, identify a precise map between the complexity’s degrees of freedom and the gravitational analogues. Alternatively, one could use bulk symplectic forms to relate boundary quantities to circuit complexity [36].
@@
Complexio: Most of the holographic-complexity studies we just discussed were performed in the context of the AdS/CFT correspondence. Why have recent complexity trends focused on de Sitter spacetime? I thought that de Sitter spacetime’s holographic properties were less understood than AdS spacetime’s.
@@
Sofia: You are right; de Sitter holography is poorly understood. Nevertheless, it holds promise for understanding quantum gravity in expanding universes similar to ours. After all, we do not live in AdS spacetime. If a holographic complexity conjecture for de Sitter spacetime is correct, it could elucidate the spacetime’s quantum dual.
@@
Complexio: You mentioned earlier that the definition of quantum state complexity has ambiguities, depending on a gate set and a reference state. Does any guiding principle direct these choices?
@@
Sofia: Yes: certain operations may be easier to implement, and certain states may be easier to prepare, than others in experiments. I can illustrate with two examples. First, a superconducting qubit’s native Hamiltonian is conventionally set proportional to $\sigma_{z}$. Consequently, adding a $\sigma_{z}$ to the native Hamiltonian costs more control resources than adding a $\sigma_{x}$ or a $\sigma_{y}$ [37]. We should penalize difficult-to-implement generators. Distributed computing offers a second example of difficult operations: imagine performing a computation using multiple small quantum computers. The information processing remains highly coherent if one computer performs most of the computations. Any nonlocal operator, which acts on multiple nodes, deserves a large complexity penalty.
@@ 4 Definition and time evolution of quantum complexity
A quantum circuit consists of gates. A gate is an elementary unitary operation that acts simultaneously on some (typically small) number of qubits (or other qudits—$d$-level quantum systems). One can implement every $n$-qubit unitary arbitrarily precisely, by selecting gates from a universal set [10]. Quantum complexity is heuristically defined as the minimum number of gates needed to perform a certain task. Examples include mapping an initial state to a final one. Experimentally implementing a gate costs a finite amount of time. Therefore, some quantum-complexity measures reflect the duration of the optimal algorithm for performing a task.
In this section, we present the characteristic features of the quantum complexity’s time evolution under a chaotic Hamiltonian. They can be summarized as follows:
  • 1.
@@
One can predict the linear growth and switchback effect from simple counting arguments in a circuit model, Sec. 4.1 shows. Section 4.2 discusses the complexity’s saturation in the context of quantum many-body equilibration. In Sec. 4.3, we introduce random circuits as simple models for chaotic dynamics. In addition, random quantum circuits mimic the properties of the Haar measure with increasing depth. This phenomenon can be exploited to make many of the heuristic arguments about the growth of complexity rigorous for random quantum circuits. In Sec. 4.4 and Sec. 4.5 we present two recent rigorous arguments that prove a linear growth of quantum complexity for an exponentially long time. In Sec. 4.6 we revisit the problem of proving linear growth of quantum circuit complexity for exponentially long time for the time-evolution of time-independent local Hamiltonians from the perspective of computational complexity theory. This perspective will show that significant progress in theoretical computer science is necessary to even prove superpolynomial lower bounds on the quantum complexity in this setting.
@@ 4.1 Time evolution of complexity, according to counting arguments about quantum circuits
Consider a fast-scrambling quantum system evolving under a unitary $U(t)\coloneqq e^{-iHt}$. How does this unitary’s complexity evolve in time? We answer this question with a simple counting argument [38, 39, 1, 2].
To show how, we consider a system of $K\gg 1$ qubits. [^3] We mention qubit for concreteness, but the argument extends to higher-dimensional qudits. Suppose that the Hamiltonian $H$ is fast-scrambling: information about initially localized perturbations spreads across the system in many-body entanglement quickly (in a time logarithmic in the system size [40, 41]). Also, suppose that $H$ is 2-local, effecting just 2-body interactions. [^4] The 2 is not essential; only the interactions’ $k$-locality is, for some $k\ll K$. By discretizing time, we can regard the evolution as formed from two-qubit gates. At each time step, all the qubits are paired, and each pair undergoes a gate (Fig. 2).
@@
What is the minimal number of gates needed to simulate evolution under $H$ for some time $t$? If $t$ is, at most, exponentially large in $K$, the optimal circuit is believed to follow from discretizing the evolution under $H$. In other words, the discretized time evolution circuit likely does not contain gates that cancel each other, since the system is a fast scrambler. Of course, such cancellations would prevent the circuit from being the shortest. Hence the complexity of a fast scrambler’s $U(t)$ evolves linearly in the number $n$ of circuit layers:
@@
This is easily turned into an equation for the complexity evolution as a function of time, since $n=t/\ell$ with $\ell$ the characteristic time scale for the application of a layer in the circuit  [^5] In Sec. 4.5 we introduce $T$ for a similar quantity: the number of layers in a 1D random quantum circuit in a brickwork layout.. A precise mathematical derivation of Eq. (6) can be made in the context of random circuits (Sec. 4.4).
For how long does the unitary’s complexity grow linearly? Another counting argument answers this question, as shown in [10, Sec. 4.5.4] and [42, Sec. 8]. Until now, we have reasoned about digitally simulating evolution under the 2-local $H$. Consider, instead, an arbitrary finite, universal set of 2-qubit gates. Consider the set of all depth-$n$ circuits formed from those gates. Each circuit implements some unitary. The complexity of $U(t)$ quits growing linearly once the aforementioned circuits have explored most of the volume of the group $\mathrm{SU}(2^{K})$. [^6] This claim and the ones to follow are based on populating the group manifold with $\delta$ size balls, where $\delta$ is the tolerance. The following claims therefore hold when studying approximate circuit complexity $C_{\delta}$, see section 2. The exact derivation is reviewed in [2]. This exploration is achieved by a time $t\sim\mathcal{O}(2^{2K})$. At this time, the complexity reaches a maximal value $\mathcal{C}_{\text{max}}\sim\mathcal{O}(2^{2K})$. The complexity is expected to then fluctuate around its maximal value. After a time $\sim 2^{2^{K}}$, the complexity is expected to return to near-minimal value, in a Poincaré recurrence. Figure 3 depicts these expected features of the complexity’s evolution.
@@
In the previous two paragraphs, we supposed that gates would not cancel each other. If a circuit does contain gates that cancel, we call the cancellation a shortcut: the circuit implements the same unitary as a shorter circuit that lacks the cancelled gates. The no-shortcut assumption also enables us to answer the question how does complexity react to a perturbation?
@@
Using this precursor, we test how the system would behave if the perturbation $W$ acted an amount $t$ of time earlier. What is the complexity of $W(t)$? If $W$ is the identity operator, then the precursor’s complexity vanishes always: gates used to implement $U(t)$ always cancel gates used to implement $U(-t)$.
Gates cancel less if $W$ is nontrivial. Consider Fig. 4, a circuit that implements the precursor $W(t)$. The gates leftward of the $W$ represent the optimal construction of $U(t)$. That construction’s first gate does not commute with the perturbation $W$. Therefore, gate 1 contributes to the complexity of $W(t)$. In contrast, gate 2 commutes with the perturbation. Therefore, the gate cancels between the circuit implementations of $U(t)$ and $U(-t)$. Hence gate 2 does not contribute to the complexity of $W(t)$. Like gate 2, gate 3 commutes $W$. However, gate 3 does not (generally) commute with gate 1, since both act on the same qubits. Therefore, gate 3 likely contributes to the complexity of $W(t)$.
If a perturbation has influenced a given qubit (if the qubit is in the perturbation’s light cone) by a given time, we call that qubit infected. The number of infected qubits—the qubit epidemic—grows with the circuit depth. Once all qubits are infected, we expect gates not to cancel any longer. The epidemic crests at a time that we call the scrambling time, $t_{*}$. After this time, the complexity grows at twice the rate specified in Eq. (6).
@@
Initially, the complexity grows exponentially, as $e^{t/\ell}$. Yet the complexity is small throughout this stage, due to the $e^{-t_{*}/\ell}$ and the scrambling time’s large size. Once $t$ is large enough, we say that the perturbation has scrambled. Afterward, the complexity grows linearly, at twice the rate at which the unitary’s complexity grows [because $W(t)$ contains two unitaries]. Figure 5 illustrates the delay in the complexity’s linear growth—the switchback effect—encoded in the $-t_{*}$ in the second line of Eq. (11). Holographic complexity proposals reproduce the switchback effect and the complexity’s initial exponential growth [see Sec. 7.2, particularly the discussion around Eq. (112)]. This reproduction originally motivated the complexity conjectures discussed in Sec. 7. It is an interesting question for future research to investigate whether more formal arguments on random circuits can be used to reproduce the switchback effect. The discussion above focused on a simple perturbation of one qubit at one instant, as well as on 2-qubit gates. These assumptions can be relaxed [38, 39, 1, 2, 19, 43].

### v2

@@ 3 Dialogue
Sofia is leading a discussion about complexity. Complexio is an audience member who seeks to clarify ideas in the literature, sometimes with technical questions. Sagredo mediates.
@@
Sagredo: Finally, we resolved to meet today to clarify the past decade’s developments concerning complexity.
@@
Sofia: Complexity has diverse applications. One can use it to characterize the dynamics of a quantum system and thermalization, solve optimization problems, capture chaotic properties, describe quantum information scrambling…Shall I continue?
Sagredo: These applications sound interesting, but complexity as a concept seems intangible. In most cases, I would think that it is difficult or impossible to compute.
Sofia: Fair objection. Nevertheless, it is possible to make arguments that reveal how a state’s quantum complexity changes under generic state evolution. Also, bounding complexity is easier than calculating it, and bounds suffice for many purposes.
Sagredo: Sofia, even finding bounds—other than the simple bounds known already—seems difficult. How can we make progress?
Sofia: At least two approaches can improve current bounds on complexity. One approach centers on random unitary circuits, which are a toy model for chaotic dynamics. Invoking random unitary circuits, we can apply differential topology and algebraic geometry to manifolds formed by unitary transformations. Also, when considering random unitary circuits, we can use unitary designs—probability distributions that cannot be distinguished from the Haar measure in polynomial time.
The second approach centers on Nielsen’s notion of complexity. Suppose we want to perform some unitary transformation on $n$ qubits. Many circuits implement that transformation. Now observe that the unitary group $\mathrm{SU}(2^{n})$, equipped with a continuous norm, forms a differential manifold. Nielsen considers geodesic paths on this manifold from the identity to the transformation of interest. The geodesic is computed with respect to a “complexity metric” that specifies that distances are small in “simple” directions on the manifold and large in “complex” directions. Here “simple” can be defined in whatever way is appropriate to the problem at hand – for example, we might say that an operation is “simple” if it acts on less than $k$ qubits. The geodesic length in the complexity metric bounds the quantum complexity of the unitary transformation [15, 19]. This bound can be useful, but may not be tight.
@@
Sofia: Certainly! Kolmogorov complexity is the length of the shortest computer program that produces a target output [20]. Another notion, Krylov complexity, quantifies operator growth—an operator’s spread across the space of operators during evolution in the Heisenberg picture [21]. Similarly, spread complexity quantifies the expansion of states across the Hilbert space during time evolution [22].
@@
Sofia: Both have interesting physical applications and are relatively simple to calculate. Krylov complexity was introduced to distinguish chaotic from integrable dynamics of finite-size quantum many-body systems that lack semiclassical limits [21]. Spread complexity relates to spectral form factors, which characterize the energy spectra of many-body quantum-systems [22], and provides tools for quantifying late time dynamics. To calculate the Krylov or spread complexity, we must tridiagonalize the Hamiltonian and there are efficient numerical [21, 22] and analytical [23, 24] techniques for doing so. Therefore, one can calculate these complexities relatively easily.
@@
Sagredo: I have heard that the holographic duality is an equivalence between gravitational physics in a bulk geometry and a quantum theory defined on its boundary. Holographic duals often simplify difficult quantum problems. Does quantum complexity have a holographic interpretation?
Sofia: We hope so! Quantum complexity has been proposed to have an Einstein-Rosen bridge’s (ERB’s) volume as a holographic dual. An ERB is a spacetime region that connects a maximally extended Schwarzschild black hole’s two boundaries. In the holographic context, the black hole is in an asymptotically Anti-de Sitter (AdS) spacetime and is dual to the thermofield-double state of a Conformal Field Theory (CFT) on the spacetime boundary. The ERB’s volume was proposed to be a holographic dual to the boundary state’s quantum complexity [6, 25]. More-complicated geometric quantities were conjectured to be dual to quantum complexity, too [26, 27, 28, 29, 30].
@@
Sofia: Yes. The geometric quantities behave similarly to the complexity of a quantum state evolving under random dynamics. One important behavior is linear growth until late times. This is achieved by the proposed dual geometric quantities in a semiclassical gravitational setting. Another behavior is the switchback effect, in which a perturbation delays growth.
More evidence emerges from a class of two-dimensional bulk models. There, a notion of wormhole length was formulated nonperturbatively. This length saturates at times exponentially long in the black hole’s entropy [31] (unlike the classical theory where the length grows forever). This late time behavior mirrors the expected saturation of complexity at times exponentially long in the system size [4]. Moreover, in [32, 33, 34, 35], the authors show that a state’s spread complexity precisely matches the dual wormhole’s classical volume at early times while also saturating at late times, at least in a two-dimensional gravitational setting. What is more, in general dimensions and for any theory of gravity, the saturation of the wormhole’s size at exponential values must be enforced by the finiteness of the black hole’s entropy, as explained in [36, 37].
@@
Sofia: The map I just mentioned—between a wormhole’s volume and spread complexity—hints that spread complexity is appropriate. Yet Refs. [38, 39, 40, 41, 42, 43, 44, 45, 46] point to other notions of complexity. We are not yet certain.
@@
Complexio: The community sounds awfully uncertain about the correct notion of complexity in holographic contexts. Why bother working on the topic?
@@
Sofia: First, the gravitational and quantum settings would both merit analysis even in the absence of any duality. We want to understand the structure of the interiors of black holes and to optimize quantum circuits. The notion of complexity has facilitated these goals. Second, we do hope to replace the many conjectures with a smaller number of theorems.
@@
Complexio: The community seems to me to be far from theorems. To support a bulk-boundary map for Nielsen’s complexity, people constructed toy examples involving free QFTs. Yet free QFTs are not known to be dual to any gravitational system. Rather, certain strongly coupled QFTs are. So how can the toy examples support the map?
@@
Sofia: We expect some qualitative properties of QFTs not to depend on the coupling strength. An example is the structure of the ultraviolet divergences in the notion of complexity used in a QFT setting. The UV divergence structure of the entanglement entropy offers another example.
@@
Sagredo: Indeed, entanglement entropy provides a striking example of a holographic mapping that has passed important consistency checks. The dual to a boundary QFT state’s entanglement entropy is the area of a minimal surface in the geometry dual to that state [5]. The dual objects have been identified clearly in this example for various states.
@@
Complexio: But I heard of a paper with the title “Complexity=Anything” which seemed to imply that different notions of complexity could match each of infinitely many gravitational quantities [29, 30, 47]. Every such quantity exhibits linear growth and the switchback effect. But then, I can pick any conjecture I like, and all of them define a geometrical dual to quantum complexity. Does this multiplicity make any sense? Is the cornucopia of such holographic conjectures useful in any way?
@@
Sofia: Well, actually, as we discussed, quantum state complexity depends on a gate set and a reference state. We are free to choose each. This freedom mirrors the ambiguity in the complexity’s gravitational dual.
Still, I understand the concern – the catchphrase “Complexity = Anything” makes it sound like there is no concrete statement at all. But as we discussed earlier, there is a precise relation between spread complexity and wormhole length at least in the case of two-dimensional gravity. In terms of the “Complexity = Anything” idea, it would help to show that that the ambiguities in alternative definitions of complexity, at least within a single framework like circuit or spread complexity, precisely match the corresponding ambiguities on the gravitational side. It could also be that some symmetries or other desirable properties could restrict the appropriate gravitational quantities to one geometric object, and in this case we would need a principle to select the corresponding notion of complexity. One approach, tried in [48], was to use bulk symplectic forms to relate boundary quantities to circuit complexity.
@@
Complexio: Most of the holographic complexity studies we just discussed were performed in the context of the AdS/CFT correspondence. Why have recent works focused on de Sitter spacetime? I thought that de Sitter spacetime’s holographic properties were less understood than AdS spacetime’s.
@@
Sofia: You are right; de Sitter holography is poorly understood. Nevertheless, it holds promise for understanding quantum gravity in expanding universes similar to ours. After all, we do not live in AdS spacetime. If a holographic complexity conjecture for de Sitter spacetime were correct, it could elucidate the quantum dial of such spacetimes.
@@
Complexio: You mentioned earlier that the definition of quantum state complexity has ambiguities, and depends on a choice of gate set and a reference state. Does any guiding principle direct these choices?
@@
Sofia: Yes: certain operations may be easier to implement, and certain states may be easier to prepare, than others in experiments. I can illustrate with an example. We should penalize difficult-to-implement generators. Distributed computing offers an example of difficult operations: imagine performing a computation using multiple small quantum computers, or on widely separated qubits. It is difficult to maintain quantum coherence across such separations. It is easiest to keep the information processing coherent if one computer performs most of the computations on local qubits. Any nonlocal operator, which acts on multiple nodes, deserves a large complexity penalty.
In the alternative Krylov or spread complexity approaches, we can start with any operator or state. But [22] proved that, for a given initial state, these approaches yield a unique notion of complexity by minimizing the spread of the operator or state over all possible bases describing the evolving system. In the case of spread complexity, different initial states (e.g., random states, thermal states, states localized on some eigenspace) can be used to probe different dynamical properties of the system.
@@ 4 Definition and time evolution of quantum complexity
A quantum circuit consists of gates. A gate is an elementary unitary operation that acts simultaneously on some (typically small) number of qubits (or other qudits—$d$-level quantum systems). Every $n$-qubit unitary operation can be implemented arbitrarily precisely, by selecting an appropriate combination of gates from a universal set [15]. Quantum complexity is heuristically defined as the minimum number of gates needed to perform a certain task. Examples include mapping an initial state to a final one. Experimentally implementing a gate costs a finite amount of time. Therefore, some quantum-complexity measures reflect the duration of the optimal algorithm for performing a task.
In this section, we present the characteristic features of the time evolution of quantum complexity under the action of a chaotic Hamiltonian. These features can be summarized as follows:
@@
Sec. 4.1 shows that one can predict the linear growth and switchback effect from simple counting arguments in a circuit model. Section 4.2 discusses the saturation of complexity in the context of quantum many-body equilibration. In Sec. 4.3, we introduce random circuits as simple models for chaotic dynamics. With increasing depth, such circuits mimic the properties of unitary transformations selected from the Haar measure. This phenomenon can be exploited to make many of the heuristic arguments about the growth of complexity rigorous for random quantum circuits. In Sec. 4.4 and Sec. 4.5, we present two recent rigorous arguments that prove linear growth of quantum complexity for an exponentially long time. In Sec. 4.6 we revisit the problem of proving linear growth of quantum circuit complexity for exponentially long time for the time-evolution of time-independent local Hamiltonians from the perspective of computational complexity theory. This perspective will show that significant progress in theoretical computer science is necessary to even prove superpolynomial lower bounds on quantum complexity in this setting.
@@ 4.1 Time evolution of complexity, according to counting arguments about quantum circuits
Consider a fast-scrambling quantum system evolving under a unitary transformation $U(t)\coloneqq e^{-iHt}$. How does this unitary’s complexity evolve in time? We answer this question with a simple counting argument [49, 50, 1, 2].
To show how, consider a system of $K\gg 1$ qubits. [^2] We use qubits for concreteness; the argument extends to higher-dimensional qudits. Suppose that the Hamiltonian $H$ is fast-scrambling: information about initially localized perturbations spreads across the system in many-body entanglement quickly (in a time logarithmic in the system size [51, 52]). Also, suppose that $H$ is 2-local, effecting just 2-body interactions. [^3] We just need the interaction to be $k$-local, for some $k\ll K$. We are taking $k=2$ for concreteness. By discretizing time, we can regard the evolution as formed from two-qubit gates. At each time step, all the qubits are paired, and each pair is transformed by the action of a gate (Fig. 2).
@@
What is the minimal number of gates needed to simulate evolution under $H$ for some time $t$? If $t$ is, at most, exponentially large in $K$, the optimal circuit is believed to follow from discretizing the evolution under $H$. In other words, the discretized time evolution circuit likely does not contain gates that cancel each other, since the system is a fast scrambler. Of course, such cancellations would prevent the circuit derived from the Hamiltonian from being the shortest. Hence the complexity of a fast scrambler’s $U(t)$ should evolve linearly in the number $n$ of circuit layers:
@@
This is easily turned into an equation for the complexity evolution as a function of time, since $n=t/\ell$ with $\ell$ the characteristic time scale for the application of a layer in the circuit. [^4] In Sec. 4.5 we introduce $T$ for a similar quantity: the number of layers in a 1D random quantum circuit in a brickwork layout. A precise mathematical derivation of Eq. (6) can be made in the context of random circuits (Sec. 4.4).
For how long does the unitary’s complexity grow linearly? Another counting argument answers this question, as shown in [15, Sec. 4.5.4] and [53, Sec. 8]. So far, we have reasoned about digitally simulating evolution under a 2-local Hamiltonian. Consider, instead, an arbitrary, finite, universal set of 2-qubit gates. Consider the set of all depth-$n$ circuits formed from those gates. Each circuit implements some unitary transformation. The complexity of $U(t)$ stops growing linearly once the aforementioned circuits have explored most of the volume of the group $\mathrm{SU}(2^{K})$. [^5] This claim and the ones to follow are based on populating the group manifold with $\delta$ size balls, where $\delta$ is the tolerance. The following claims therefore hold when studying approximate circuit complexity $C_{\delta}$, see section 2. The exact derivation is reviewed in [2]. This exploration is achieved by a time $t\sim\mathcal{O}(2^{2K})$. At this time, the complexity reaches a maximal value $\mathcal{C}_{\text{max}}\sim\mathcal{O}(2^{2K})$. The complexity is expected to then fluctuate around its maximal value. After a time $\sim 2^{2^{K}}$, the complexity is expected to return to near-minimal value, in a Poincaré recurrence. Figure 3 depicts these expected features of the evolution of complexity.
@@
In the previous two paragraphs, we supposed that gates would not cancel each other’s effects. If a circuit does contain gates that cancel, we call the cancellation a shortcut: the circuit implements the same unitary as a shorter circuit that lacks the cancelled gates. The no-shortcut assumption also enables us to answer the question how does complexity react to a perturbation?
@@
Using this precursor, we can test how the system would behave if the perturbation $W$ acted an amount $t$ of time earlier. What is the complexity of $W(t)$? If $W$ is the identity operator, then the precursor’s complexity vanishes always: gates used to implement $U(t)$ always cancel gates used to implement $U(-t)$.
Gates cancel less if $W$ is nontrivial. Consider Fig. 4, a circuit that implements the precursor $W(t)$. The gates leftward of $W$ represent the optimal construction of $U(t)$. That construction’s first gate does not commute with the perturbation $W$. Therefore, gate 1 contributes to the complexity of $W(t)$. In contrast, gate 2 commutes with the perturbation. Therefore, the gate cancels between the circuit implementations of $U(t)$ and $U(-t)$. Hence gate 2 does not contribute to the complexity of $W(t)$. Like gate 2, gate 3 commutes with $W$. However, gate 3 does not (generally) commute with gate 1, since both act on the same qubits. Therefore, gate 3 likely contributes to the complexity of $W(t)$.
If a perturbation has influenced a given qubit (if the qubit is in the perturbation’s light cone) by a given time, we call that qubit infected. The number of infected qubits—the qubit epidemic—grows with the circuit depth. Once all qubits are infected, we expect gates not to cancel any longer. The epidemic crests at a time that we call the scrambling time, $t_{*}$. After this time, as we will see, the complexity grows at twice the rate specified in Eq. (6).
@@
Initially, the complexity grows exponentially, as $e^{t/\ell}$. Yet the complexity is small throughout this stage, due to the $e^{-t_{*}/\ell}$ and the scrambling time’s large size. Once $t$ is large enough, we say that the perturbation has scrambled. Afterward, the complexity grows linearly, at twice the rate at which the unitary’s complexity grows, because $W(t)$ contains two unitaries. Figure 5 illustrates the delay in the complexity’s linear growth—the switchback effect—encoded in the $-t_{*}$ in the second line of Eq. (11). Holographic complexity proposals reproduce the switchback effect and the initial exponential growth of complexity (see Sec. 7.2, particularly the discussion around Eq. (123)). This fact originally motivated the complexity conjectures discussed in Sec. 7. An interesting question for future research is to investigate whether more formal arguments with random circuits can be used to reproduce the switchback effect. The discussion above focused on a simple perturbation of one qubit at one instant, as well as on 2-qubit gates. These assumptions can be relaxed [49, 50, 1, 2, 25, 54].